OpenStudy (anonymous):

Word Problem. An inheritance of $800,000 is to be divided among Avery, Connor and Olivia in the following matter: Olivia is to recieve 3/4 of what Connor gets, while Avery gets 1/4 of what Connor gets. How much does each recieves? Show me steps by steps please

OpenStudy (anonymous):

okay if you don't mind can i say 800 instead of 800 000 just for simplicity sake ? we can add it on after?

OpenStudy (anonymous):

i'm not here to DO your homework if you want help, you'll have to work through it with me

OpenStudy (anonymous):

okay I work with you.

OpenStudy (anonymous):

okay so at first we're going to solve for just one of the siblings, in terms of one variable. Looks like the easiest will be Connor so we will let Connor = C

OpenStudy (anonymous):

O for Olivia

OpenStudy (anonymous):

so if Olivia gets (3/4) of what Connor gets then we know that Olvia = (3/4)*C

OpenStudy (anonymous):

so knowing that what does Avery get? in terms of C

OpenStudy (anonymous):

uh huh. Avery get 1/4 of what Connor get

OpenStudy (anonymous):

right, so.... that looks like (1/4)*C..... right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

okay so now the amount each sibling gets is in terms of Connor or C. so the sum of all of their values will be the inheritance right?

OpenStudy (anonymous):

Can you come up with an equation for me?

OpenStudy (anonymous):

so you said 1/4+3/4+x=800,000

OpenStudy (anonymous):

almost! except you can't forget to keep the x's tagged on the terms. I suggest we keep doing it in terms of C, since it will be clearer, there's no need to switch to x now.

OpenStudy (anonymous):

so: (1/4)C+(3/4)C+C=800,000

OpenStudy (anonymous):

ok. so you add up the fraction right?

OpenStudy (anonymous):

right and you will see right away it will simplify nicely!

OpenStudy (anonymous):

that meant that 4/4 is equal 1 then you subtract 1 from 800,000 right?

OpenStudy (anonymous):

you keep forgetting the C!! but the 4/4=1 is right.. so now you have (4/4)C+C=800,000 or.. 1C+1C=800,000

OpenStudy (anonymous):

so there will be no subtraction in this question

OpenStudy (anonymous):

2C=800,000 then you divide by 2 right?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

the Connor get 400,000. then Olivia get 300,000 because you multiply 3/4 to 400,000. Then Avery get 100,000 because I multiply 1/4 to 400.000 right?

OpenStudy (anonymous):

nailed it.

OpenStudy (anonymous):

oh I get it.

OpenStudy (anonymous):

I just hate word problems

OpenStudy (anonymous):

you can check this by adding 400 000 + 300 000 + 100 000 = 800 000

OpenStudy (anonymous):

remember that's always an important step.